# Improving Group Robustness on Spurious Correlation Requires Preciser Group Inference

### v1

@@ B.2.1 Dataset Details
CMNIST[2] is a noisy digit recognition task. The binary feature (green and red), referred to as color, serves as a spurious feature, while the binary feature (digit contours) acts as the invariant feature. The CMNIST dataset involves two classes, where class 0 corresponds to the original digits (0,1,2,3,4), and class 1 represents digits (5,6,7,8,9). Following the approach recommended in [50], we construct a training set with a sample size of 30,000. In class 0, the ratio of red to green samples is set at 8:2, while in class 1, it is set at 2:8. For the validation set consisting of 10,000 samples, the proportion of green to red samples is equal at 1:1 for all classes. The test set, containing 20,000 samples, features a proportion of green to red samples at 1:9 in class 0 and 9:1 in class 1. Additionally, label flipping is applied with a probability of 0.25. In our experiments, we utilized the validation set as the comparison data and employed an unlabeled validation set to simulate scenarios where comparison data is unavailable. The four groups of CMNIST is $(g_{1},g_{2},g_{3},g_{4})=(\{0,green\},\{0,red\},\{1,green\},\{1,red\})$ and the group distribution of the training data is $(g_{1},g_{2},g_{3},g_{4})=(0.1,0.4,0.4,0.1)$, while the group distribution of the comparison data is $(g_{1},g_{2},g_{3},g_{4})=(0.26,0.25,0.25,0.24)$.
@@
Waterbirds aims to classify bird images as either waterbirds or landbirds, with each bird image falsely associated with either a water or land background. Waterbirds is a synthetic dataset where each image is generated by combining bird images sampled from the CUB dataset [47] with backgrounds selected from the Places dataset [52]. We directly load the Waterbirds dataset using the Wilds library in PyTorch [23]. The dataset consists of a total of 4,795 training samples, with only 56 samples labeled as waterbirds on land and 184 samples labeled as landbirds on water. The remaining training data includes 3,498 samples from landbirds on land and 1,057 samples from waterbirds on water. We still directly use the validation set as comparison data. The waterbirds dataset can be divided into four groups, namely $(g_{1},g_{2},g_{3},g_{4})=(\{landbird,land\},\{landbird,water\},\{waterbird,land\},\{waterbird,water\})$. The class distribution of the training data is $(g_{1},g_{2},g_{3},g_{4})=(0.73,0.04,0.10,0.13)$, while the class distribution of the comparison data is $(g_{1},g_{2},g_{3},g_{4})=(0.56,0.22,0.16,0.16)$.
@@
CelebA [32, 40] is a hair-color prediction task, similar to the study conducted by [50], and follows the data preprocessing procedure outlined in [40]. Given facial images of celebrities as input, the task is to identify their hair color as either blond or non-blond. This labeling is spuriously correlated with gender, which can be either male or female. In the training set, there are 71,629 instances (44%) of females with non-blond hair, 66,874 instances (41%) of non-blond males, 22,880 instances (14%) of blond females, and 1,387 instances (1%) of blond males. In the validation set, there are 8535 instances (43%) of females with non-blond hair, 8276 instances (42%) of non-blond males, 2874 instances (14%) of blond females, and 182 instances (1%) of blond males. The validation set is still regarded as the comparison data to infer group labels. We found that the group distribution of the training data and the comparison data in the CelebA dataset are very similar, which poses a challenge for GIC inference.
@@ B.2.2 Baseline Details
Spurious correlations mitigation with group labels. GroupDRO [40] is a well-established method that enhances worst-group accuracy using group labels. Unlike standard ERM, which minimizes the average loss across all groups, GroupDRO partitions the data based on prior knowledge of groups and minimizes the worst-case loss over groups in the training data, thereby improving worst group accuracy. DFR [22] balances the training set by leveraging the Subsample strategy. This involves retaining all data from the smallest group and Subsample data from the other groups to equalize group sizes, followed by retraining the ERM model using this balanced dataset. LISA [50] addresses spurious correlation between the spurious attribute and true label through Mixup. It employs two mixup strategies: Intra-label, which interpolates samples with the same label but different spurious attributes, and intra-domain, which interpolates samples with the same spurious attributes but different true labels. Through these mixup techniques, LISA achieves a significant improvement in the worst-group accuracy. It is important to emphasize that LISA utilizes different Mixup techniques for various experimental datasets. Specifically, for CMNIST and Waterbirds, the classic mixup (i.e., linear interpolation between two images) is considered. In the case of CelebA, CutMix, which involves replacing removed regions with a patch from another image, is employed. Meanwhile, for CivilComments, LISA resorts to using Manifold Mix. For all four datasets in the experiments, GIC adopts the same Mixup strategies as LISA.
@@
Spurious correlations mitigation without group labels. JTT [31] initially treats misclassified points by a trained ERM as errors, which are then upsampled for improving model robustness. Similarly, CnC [51] uses the outputs of an ERM model to identify samples with the same class but different spurious features, and trains a robust model with contrastive learning. EIIL [12] learns spurious attribute (group) labels using the EI method, which maximizes the loss of GroupDRO, and utilizes GroupDRO to learn invariant features. However, its performance is unstable and relatively poor in scenarios where learning spurious correlations is challenging [30]. SSA [35] leverages semi-supervised learning with available group labels for some samples to train a spurious attribute predictor. ZIN [30] improves group inference with auxiliary information such as timestamps for time-series data or meta-annotations for images, while DISC [48] constructs a concept bank with potential spurious attributes. However, acquiring such auxiliary information (human priors or partial group labels) may also pose difficulties, limiting the general applicability of these methods.
@@
Downstream invariant learning methods. We considere four downstream invariant learning algorithms. Mixup [50] selectively mixes samples with the same label but different spurious features or samples with different labels but the same spurious feature, to achieve invariance. Subsample [22] constructs balanced data by keeping the smallest sample size group and subsampling other groups to the same quantity. Upsample [31] oversamples samples in the misclassified sets using ERM, emphasizing difficult-to-predict groups, as well as the classic GrouDRO [40] algorithm.
@@ B.2.3 Training Details
Stage 1: Extracting feature representations. First, we introduce the feature extractors $\Phi(\cdot)$ used for each dataset. We follow previous works [22, 31, 51, 50] and consider architectures including LeNet-5 CNN (CMNIST), ResNet-50 (Waterbirds and CelebA), and BERT (CivilComments) as our feature extractors. We detail the feature extractors architecture and hyperparameters for each dataset below:
@@
The CelebA and CivilComments datasets directly use pretrained models as feature extractors without any fine-tuning, and CMNSIT trains the extractor from scratch. It is worth noting that the Waterbirds dataset utilizes pretrained weights from ImageNet and then fine-tunes them on the Waterbirds dataset. We adopt this strategy because previous work [22] has observed that initializing the feature extractor with ImageNet trained weights and fine-tuning on the dataset can significantly improve the performance of feature extraction on the Waterbirds dataset. However, such improvements are not significant on datasets with larger sample sizes, such as CelebA.
@@
Stage 3: Invariant learning. After obtaining the inferred group labels using GIC, the final stage is the invariant learning stage. In the third stage, we train the robust model using the inferred group labels from the second stage for invariant learning. In our experiments, we consider four invariant learning algorithms: Mixup, GroupDRO, Subsample, and Upsample. During this stage, we use the parameters and model architecture corresponding to each specific invariant learning algorithm. For Mixup, we directly use the parameters from the original LISA paper [50]. We only need to replace the spurious class labels of the original training set with the labels predicted by GIC. For GroupDRO and Subsample, we follow the parameters implemented in EIIL [12], CnC [51], and DFR [22]. As for Upsample, we refer to the parameters selected by JTT [31].
@@ Appendix C More Results
In this section, we provide additional experimental results as a supplement to Section 4. Firstly, we fully display the average group accuracy and the worst-group accuracy as outlined in Section 4.4. Subsequently, we showcase the results of ablation experiments, illustrating the performance enhancement of GIC with the adoption of early stopping. Then, we present further error cases from the Waterbirds and CelebA datasets, underscoring the prevalent manifestation of semantic consistency in GIC.Following that, we discuss the various sources of comparison data and the impact of their sample sizes as mentioned in Section 3.4, to elucidate the effect of distinct origins and quantities of comparison data on the performance of GIC. Finally, we highlight the performance implications of the distribution discrepancy between the comparison data and training data on GIC, augmenting the insights in Section 5.
@@ C.1 Complete Results of group inference method comparison
In this section, we present the experimental results that were not fully shown in section 4.4 due to space limitations. These results include the average group accuracy, where GIC outperforms baselines not only in the worst-performing group but also achieves higher accuracy in the average group. This further demonstrates that GIC has the ability to improve both the average and worst-group accuracy.
Table 3: Average and worst group accuracy comparison (%).
@@ C.2 Ablation Study
In this section, We present ablation experiments results. For each invariant learning algorithms, including Mixup, GroupDRO, Upsample, and Subsample, we followe previous research [31, 50, 51, 12],and us the group labels of the validation set and consider the worst-group accuracy to determine early stopping. Tables 4 and 5 show the average and worst-performing group performances obtained by combining GIC with different invariant learning algorithms, both with and without early stopping. We observe that the reference for early stopping does indeed improve performance, especially for the worst-group.
@@ C.3 More error cases
In Section 4.5, we illustrate the semantic consistency of GIC using the celebA dataset, where images with similar semantics are considered as belonging to the same spurious feature category. This distinguishes GIC, which relies on semantic features for recognition, from human-based oracle group labels. We observe the same phenomenon on the waterbirds dataset, as shown in Figure 7. The misclassifications of GIC can be categorized into two types: (1) Land background with typical water features, such as extensive blue regions, often leads GIC to misclassify land as water. (2) Water background with typical land features, such as abundant tree branches, ponds with lush green vegetation, or large tree reflections, frequently results in GIC classifying them as land. We believe that such consistency precisely reflects GIC’s accurate recognition of image semantics. As mentioned in the main text, such accurate recognition is crucial when it comes to invariant learning that requires leveraging different image semantics (e.g., using mixup to disrupt semantic features of spurious features). We present more examples of misclassifications by GIC from the Waterbirds and CelebA datasets in Figures 8, 9, and 10 to support the existence of semantic consistency in GIC.
@@ C.4 The construction of comparison data
In this section, we discuss the methods for constructing comparison data. The availability of comparison data that satisfies the different group distributions requirement is a key factor in implementing GIC and directly determines its usability. In Section 3.4, we proposed three methods for obtaining comparison data. In the experimental section (Section 4), we used labeled/unlabeled validation sets as labeled/unlabeled comparison data from non-training datasets. The significant improvement in worst-group accuracy indicates the effectiveness of these two methods of obtaining comparison data.
@@
To support the effectiveness of the method of constructing comparison data by non-uniform sampling from the training set, we also conducted additional experiments on CMNIST. Following a similar approach to JTT, we train an ERM model on the training set and divide the samples into an error set (misclassified samples) and a non-error set (correctly classified samples). The error set is typically composed of samples from the minority group with spurious associations [31]. We then sample an equal number of samples from both the error set and non-error set. For instance, when the sampling ratio is set to 1%, we sample 300 (1% of 30,000) samples from each set, resulting in a total of 600 samples as the comparison data. These comparison data samples are combined with the remaining 29,400 training samples for GIC training. After obtaining the spurious attribute classifier $f_{\mathrm{GIC}}$, we directly evaluate its prediction accuracy on the entire training dataset, as shown in Figure 11.
@@
Figure 11: Test Accuracy on Spurious Attributes Using Non-Uniform Sampling from the Training Dataset. The dashed line represents the accuracy of GIC’s spurious feature label inference using labeled validation data. We observed that as the sampling rate gradually increases, leading to a growth in the number of comparison data samples, GIC’s capability to infer spurious features also enhances. Remarkably, on CMNIST, when the sampling rate is around 10%, the performance of GIC based on training data for inferring spurious features aligns with that of GIC using validation data directly.
@@
Furthermore, we want to emphasize the necessity of considering unlabeled comparison data, such as sampling from the test set. In the real world, unlabeled data is often more cost-effective and easier to obtain than labeled data. By utilizing a large amount of unlabeled data as comparison data, the applicability of GIC can be greatly expanded. Our results in Figure 12 demonstrate that when GIC trains a spurious feature classifier using labeled comparison data (with sample sizes ranging from 10 to 1000), its accuracy in predicting spurious features and improving the accuracy of the worst-performing group is significantly worse than when using unlabeled but larger comparison data (with sample sizes ranging from 3000 to 10000). This experiment highlights the importance of considering unlabeled comparison data.
@@
C.5 Group Distribution Discrepancy Analysis

### v2

@@ B.2.1 Dataset Details
CMNIST(Arjovsky et al., 2019) is a noisy digit recognition task. The binary feature (green and red), referred to as color, serves as a spurious feature, while the binary feature (digit contours) acts as the invariant feature. The CMNIST dataset involves two classes, where class 0 corresponds to the original digits (0,1,2,3,4), and class 1 represents digits (5,6,7,8,9). Following the approach recommended in (Yao et al., 2022), we construct a training set with a sample size of 30,000. In class 0, the ratio of red to green samples is set at 8:2, while in class 1, it is set at 2:8. For the validation set consisting of 10,000 samples, the proportion of green to red samples is equal at 1:1 for all classes. The test set, containing 20,000 samples, features a proportion of green to red samples at 1:9 in class 0 and 9:1 in class 1. Additionally, label flipping is applied with a probability of 0.25. In our experiments, we utilized the validation set as the comparison data and employed an unlabeled validation set to simulate scenarios where comparison data is unavailable. The four groups of CMNIST is $(g_{1},g_{2},g_{3},g_{4})=(\{0,green\},\{0,red\},\{1,green\},\{1,red\})$ and the group distribution of the training data is $(g_{1},g_{2},g_{3},g_{4})=(0.1,0.4,0.4,0.1)$, while the group distribution of the comparison data is $(g_{1},g_{2},g_{3},g_{4})=(0.26,0.25,0.25,0.24)$.
@@
Waterbirds aims to classify bird images as either waterbirds or landbirds, with each bird image falsely associated with either a water or land background. Waterbirds is a synthetic dataset where each image is generated by combining bird images sampled from the CUB dataset (Wah et al., 2011) with backgrounds selected from the Places dataset (Zhou et al., 2017). We directly load the Waterbirds dataset using the Wilds library in PyTorch (Koh et al., 2021). The dataset consists of a total of 4,795 training samples, with only 56 samples labeled as waterbirds on land and 184 samples labeled as landbirds on water. The remaining training data includes 3,498 samples from landbirds on land and 1,057 samples from waterbirds on water. We still directly use the validation set as comparison data. The waterbirds dataset can be divided into four groups, namely $(g_{1},g_{2},g_{3},g_{4})=(\{landbird,land\},\{landbird,water\},\{waterbird,land\},\{waterbird,water\})$. The class distribution of the training data is $(g_{1},g_{2},g_{3},g_{4})=(0.73,0.04,0.10,0.13)$, while the class distribution of the comparison data is $(g_{1},g_{2},g_{3},g_{4})=(0.56,0.22,0.16,0.16)$.
@@
CelebA (Liu et al., 2015; Sagawa et al., 2019) is a hair-color prediction task, similar to the study conducted by (Yao et al., 2022), and follows the data preprocessing procedure outlined in (Sagawa et al., 2019). Given facial images of celebrities as input, the task is to identify their hair color as either blond or non-blond. This labeling is spuriously correlated with gender, which can be either male or female. In the training set, there are 71,629 instances (44%) of females with non-blond hair, 66,874 instances (41%) of non-blond males, 22,880 instances (14%) of blond females, and 1,387 instances (1%) of blond males. In the validation set, there are 8535 instances (43%) of females with non-blond hair, 8276 instances (42%) of non-blond males, 2874 instances (14%) of blond females, and 182 instances (1%) of blond males. The validation set is still regarded as the comparison data to infer group labels. We found that the group distribution of the training data and the comparison data in the CelebA dataset are very similar, which poses a challenge for GIC inference.
@@ B.2.2 Baseline Details
Spurious correlations mitigation with group labels. GroupDRO (Sagawa et al., 2019) is a well-established method that enhances worst-group accuracy using group labels. Unlike standard ERM, which minimizes the average loss across all groups, GroupDRO partitions the data based on prior knowledge of groups and minimizes the worst-case loss over groups in the training data, thereby improving worst-group accuracy. DFR (Kirichenko et al., 2022) balances the training set by leveraging the Subsample strategy. This involves retaining all data from the smallest group and Subsample data from the other groups to equalize group sizes, followed by retraining the ERM model using this balanced dataset. LISA (Yao et al., 2022) addresses spurious correlation between the spurious attribute and true label through Mixup. It employs two mixup strategies: Intra-label, which interpolates samples with the same label but different spurious attributes, and intra-domain, which interpolates samples with the same spurious attributes but different true labels. Through these mixup techniques, LISA achieves a significant improvement in the worst-group accuracy. It is important to emphasize that LISA utilizes different Mixup techniques for various experimental datasets. Specifically, for CMNIST and Waterbirds, the classic mixup (i.e., linear interpolation between two images) is considered. In the case of CelebA, CutMix, which involves replacing removed regions with a patch from another image, is employed. Meanwhile, for CivilComments, LISA resorts to using Manifold Mix. For all four datasets in the experiments, GIC adopts the same Mixup strategies as LISA.
@@
Spurious correlations mitigation without group labels. JTT (Liu et al., 2021) initially treats misclassified points by a trained ERM as errors, which are then upsampled for improving model robustness. Similarly, CnC (Zhang et al., 2022) uses the outputs of an ERM model to identify samples with the same class but different spurious features, and trains a robust model with contrastive learning. EIIL (Creager et al., 2021) learns spurious attribute (group) labels using the EI method, which maximizes the loss of GroupDRO, and utilizes GroupDRO to learn invariant features. However, its performance is unstable and relatively poor in scenarios where learning spurious correlations is challenging (Lin et al., 2022). SSA (Nam et al., 2022) leverages semi-supervised learning with available group labels for some samples to train a spurious attribute predictor. ZIN (Lin et al., 2022) improves group inference with auxiliary information such as timestamps for time-series data or meta-annotations for images, while DISC (Wu et al., 2023) constructs a concept bank with potential spurious attributes. However, acquiring such auxiliary information (human priors or partial group labels) may also pose difficulties, limiting the general applicability of these methods.
@@
Downstream invariant learning methods. We considere four downstream invariant learning algorithms. Mixup (Yao et al., 2022) selectively mixes samples with the same label but different spurious features or samples with different labels but the same spurious feature, to achieve invariance. Subsample (Kirichenko et al., 2022) constructs balanced data by keeping the smallest sample size group and subsampling other groups to the same quantity. Upsample (Liu et al., 2021) oversamples samples in the misclassified sets using ERM, emphasizing difficult-to-predict groups, as well as the classic GrouDRO (Sagawa et al., 2019) algorithm.
@@ B.2.3 Training Details
Stage 1: Extracting feature representations. First, we introduce the feature extractors $\Phi(\cdot)$ used for each dataset. We follow previous works (Kirichenko et al., 2022; Liu et al., 2021; Zhang et al., 2022; Yao et al., 2022) and consider architectures including LeNet-5 CNN (CMNIST), ResNet-50 (Waterbirds and CelebA), and BERT (CivilComments) as our feature extractors. We detail the feature extractors architecture and hyperparameters for each dataset below:
@@
The CelebA and CivilComments datasets directly use pretrained models as feature extractors without any fine-tuning, and CMNSIT trains the extractor from scratch. It is worth noting that the Waterbirds dataset utilizes pretrained weights from ImageNet and then fine-tunes them on the Waterbirds dataset. We adopt this strategy because previous work (Kirichenko et al., 2022) has observed that initializing the feature extractor with ImageNet trained weights and fine-tuning on the dataset can significantly improve the performance of feature extraction on the Waterbirds dataset. However, such improvements are not significant on datasets with larger sample sizes, such as CelebA.
@@
Stage 3: Invariant learning. After obtaining the inferred group labels using GIC, the final stage is the invariant learning stage. In the third stage, we train the robust model using the inferred group labels from the second stage for invariant learning. In our experiments, we consider four invariant learning algorithms: Mixup, GroupDRO, Subsample, and Upsample. During this stage, we use the parameters and model architecture corresponding to each specific invariant learning algorithm. For Mixup, we directly use the parameters from the original LISA paper (Yao et al., 2022). We only need to replace the spurious class labels of the original training set with the labels predicted by GIC. For GroupDRO and Subsample, we follow the parameters implemented in EIIL (Creager et al., 2021), CnC (Zhang et al., 2022), and DFR (Kirichenko et al., 2022). As for Upsample, we refer to the parameters selected by JTT (Liu et al., 2021).
@@ Appendix C More Results
In this section, we provide additional experimental results as a supplement to Section 4. Firstly, we fully display the average-group accuracy and the worst-group accuracy as outlined in Section 4.4. Subsequently, we showcase the results of ablation experiments, illustrating the performance enhancement of GIC with the adoption of early stopping. Then, we present further error cases from the Waterbirds and CelebA datasets, underscoring the prevalent manifestation of semantic consistency in GIC.Following that, we discuss the various sources of comparison data and the impact of their sample sizes as mentioned in Section 3.4, to elucidate the effect of distinct origins and quantities of comparison data on the performance of GIC. Finally, we highlight the performance implications of the distribution discrepancy between the comparison data and training data on GIC, augmenting the insights in Section 5.
@@ C.1 Complete Results of group inference method comparison
In this section, we present the experimental results that were not fully shown in Section 4.4 due to space limitations. These results include the average-group accuracy, where GIC outperforms baselines not only in the worst-performing group but also achieves higher accuracy in the average group. This further demonstrates that GIC has the ability to improve both the average and worst-group accuracy.
Table 3: Average and worst-group accuracy comparison (%).
@@ C.2 Ablation Study
In this section, We present ablation experiments results. For each invariant learning algorithms, including Mixup, GroupDRO, Upsample, and Subsample, we followe previous research (Liu et al., 2021; Yao et al., 2022; Zhang et al., 2022; Creager et al., 2021),and us the group labels of the validation set and consider the worst-group accuracy to determine early stopping. Tables 4 and 5 show the average and worst-performing group performances obtained by combining GIC with different invariant learning algorithms, both with and without early stopping. We observe that the reference for early stopping does indeed improve performance, especially for the worst-group.
@@ C.3 More Error Cases
In Section 4.5, we illustrate the semantic consistency of GIC using the celebA dataset, where images with similar semantics are considered as belonging to the same spurious feature category. This distinguishes GIC, which relies on semantic features for recognition, from human-based oracle group labels. We observe the same phenomenon on the waterbirds dataset, as shown in Figure 7. The misclassifications of GIC can be categorized into two types: (1) Land background with typical water features, such as extensive blue regions, often leads GIC to misclassify land as water. (2) Water background with typical land features, such as abundant tree branches, ponds with lush green vegetation, or large tree reflections, frequently results in GIC classifying them as land. We believe that such consistency precisely reflects GIC’s accurate recognition of image semantics. As mentioned in the main text, such accurate recognition is crucial when it comes to invariant learning that requires leveraging different image semantics (e.g., using mixup to disrupt semantic features of spurious features). We present more examples of misclassifications by GIC from the Waterbirds and CelebA datasets in Figure 8, 9, and 10 to support the existence of semantic consistency in GIC.
@@ C.4 The Construction of Comparison Data
In this section, we discuss the methods for constructing comparison data. In Section 3.4, we proposed three methods for obtaining comparison data. In the experimental section (Section 4), we used labeled/unlabeled validation sets as labeled/unlabeled comparison data from non-training datasets. The significant improvement in worst-group accuracy indicates the effectiveness of these two methods of obtaining comparison data.
@@
To support the effectiveness of the method of constructing comparison data by non-uniform sampling from the training set, we first conduct additional experiments on CMNIST. Following a similar approach to JTT, we train an ERM model on the training set and divide the samples into an error set (misclassified samples) and a non-error set (correctly classified samples). The error set is typically composed of samples from the minority group with spurious associations (Liu et al., 2021). We then sample an equal number of samples from both the error set and non-error set. For instance, when the sampling ratio is set to 1%, we sample 300 (1% of 30,000) samples from each set, resulting in a total of 600 samples as the comparison data. These comparison data samples are combined with the remaining 29,400 training samples for GIC training. After obtaining the spurious attribute classifier $f_{\mathrm{GIC}}$, we directly evaluate its prediction accuracy on the entire training dataset, as shown in Figure 11. We observe that as the sampling rate gradually increases, leading to a growth in the number of comparison data samples, GIC’s capability to infer spurious features also enhances. Remarkably, on CMNIST, when the sampling rate is around 10%, the performance of GIC based on training data for inferring spurious features aligns with that of GIC using validation data directly.
@@
Figure 11: Test accuracy on spurious attributes using non-uniform sampling from the training dataset. The dashed line represents the accuracy of GIC’s spurious feature label inference using labeled validation data.
We also conduct additional experiments on CMNIST, Waterbirds, and CelebA to explore the strategy of non-uniform sampling from training data as comparison data. Notably, we observe that for Waterbirds and CelebA, relying solely on training data would significantly limit the sample size of comparison data. Therefore, we combine the training and validation data as a new dataset for constructing comparison data for the Waterbirds and CelebA datasets. Specifically, we mix the training and validation data, then split them into new training and validation sets at a 50% ratio. We train on the new training set using the ERM method and infer labels on the new validation set, resulting in an error set (misclassified samples) and a non-error set (correctly classified samples). Similarly, we train the model on the validation set and identified the error set and non-error set on the training set through group inference. We then combine the error and non-error sets and sample a fixed proportion (50%) from each set to serve as comparison data, with the remainder as training data. We then apply GIC to infer group labels and learn invariant features. To further enhance performance, we implement a boosting approach by repeating the aforementioned process twice. In the second iteration, we base it on the GIC model obtained from the first round of training to infer group labels and construct the error set and non-error set.
We further report the worst-group accuracy of GIC when the comparison data is generated by non-uniform sampling from the training dataset, using Mixup as the downstream algorithm. We refer to this variant of GIC as $\mathrm{GIC}_{\mathcal{C}_{t}}$-M. The experimental results in Table 6 demonstrates that constructing comparison data by non-uniform sampling from the training dataset is an effective strategy, as the worst-group accuracy is close to that achieved by the GIC approach using labeled/unlabeled validation data as the comparison dataset, particularly on CMNIST, where the performances are almost comparable.
Table 6: Comparison of GIC’s worst-group accuracy using different sources of comparison data.
Method CMNIST Waterbirds CelebA
validation
$\mathrm{GIC}_{\mathcal{C}_{y}}$-M 72.2±$\pm\scriptstyle{0.5}$ 86.3±$\pm\scriptstyle{0.1}$ 89.4±$\pm\scriptstyle{0.2}$
$\mathrm{GIC}_{\mathcal{C}}$-M 71.7±$\pm\scriptstyle{0.3}$ 85.4±$\pm\scriptstyle{0.1}$ 89.5±$\pm\scriptstyle{0.0}$
non-uniform sampling.
$\mathrm{GIC}_{\mathcal{C}_{t}}$-M 72.2 ±$\pm\scriptstyle{0.1}$ 85.7±$\pm\scriptstyle{0.3}$ 87.5±$\pm\scriptstyle{0.7}$
C.5 The Necessity of Considering Unlabeled Datasets as Comparison Data.
@@
Additionally, we want to emphasize the necessity of considering unlabeled comparison data, such as sampling from the test set. In the real world, unlabeled data is often more cost-effective and easier to obtain than labeled data. By utilizing a large amount of unlabeled data as comparison data, the applicability of GIC can be greatly expanded. Our results in Figure 12 demonstrate that when GIC trains a spurious feature classifier using labeled comparison data (with sample sizes ranging from 10 to 1000), its accuracy in predicting spurious features and improving the accuracy of the worst-performing group is significantly worse than when using unlabeled but larger comparison data (with sample sizes ranging from 3000 to 10000). This experiment highlights the importance of considering unlabeled comparison data.
@@
C.6 Group Distribution Discrepancy Analysis